### v1

# A general approach to modeling environmental mixtures with multivariate outcomes

Abstract
An important goal of environmental health research is to assess the health risks posed by mixtures of multiple environmental exposures. In these mixtures analyses, flexible models like Bayesian kernel machine regression and multiple index models are appealing because they allow for arbitrary non-linear exposure-outcome relationships. However, this flexibility comes at the cost of low power, particularly when exposures are highly correlated and the health effects are weak, as is typical in environmental health studies. We propose an adaptive index modelling strategy that borrows strength across exposures and outcomes by exploiting similar mixture component weights and exposure-response relationships. In the special case of distributed lag models, in which exposures are measured repeatedly over time, we jointly encourage co-clustering of lag profiles and exposure-response curves to more efficiently identify critical windows of vulnerability and characterize important exposure effects. We then extend the proposed approach to the multivariate index model setting where the true index structure—the number of indices and their composition—is unknown, and introduce variable importance measures to quantify component contributions to mixture effects. Using time series data from the National Morbidity, Mortality and Air Pollution Study, we demonstrate the proposed methods by jointly modelling three mortality outcomes and two cumulative air pollution measurements with a maximum lag of 14 days.
1 Introduction
Quantifying associations between health outcomes and mixtures of environmental pollutants is a public health priority, and addressing the methodological challenges introduced by mixtures analyses has been the focus of much recent research (Davalos et al.,, 2017; Stafoggia et al.,, 2017; Hamra and Buckley,, 2018; Maitre et al.,, 2022; Yu et al.,, 2022; Tanner et al.,, 2020; Joubert et al.,, 2022). Common to mixtures studies is a lack of power—owing in part to limited sample sizes, highly correlated exposures, and small effect sizes. This is exacerbated by the fact that environmental exposures may have non-linear relationships with health outcomes, so the functional form of these relationships must also be estimated from limited data. Moreover, in studies of time-varying exposures, the timing of exposure effects is itself unknown: for example studies of prenatal exposure to metals attempt to identify windows of time in which children are most susceptible to exposure effects (Barr Jr et al.,, 2000).
Many methods have been proposed to solve one or more of these issues. To handle high correlation among mixture components in a linear model, Dirichlet Process (DP) priors have been placed on coefficients (Herring,, 2010; Dunson et al.,, 2008; LaLonde et al.,, 2020), among other latent variable approaches (Zavez et al.,, 2023). Non-linear relationships have been captured by semi-parametric regression using splines (Antonelli et al.,, 2020; Bai et al.,, 2022) and Gaussian processes (Bobb et al.,, 2015; Ferrari and Dunson,, 2019), as well as by single (Wang et al.,, 2020) and multiple index models (McGee et al.,, 2021, 2024). When exposures are measured repeatedly over time, distributed lag models (DLMs) have smoothed association parameters over time and characterized time-lag profiles that identify periods in which risks are highest (Gasparrini et al.,, 2010; Gasparrini,, 2014; Gasparrini et al.,, 2017; Wilson et al.,, 2017, 2022; Antonelli et al.,, 2024). While most of these approaches are intended to address multiple exposures, multiple time points, or both, none additionally address the complexity stemming from having multiple outcomes. It is common in environmental research to be interested in a variety of health outcomes, and analyzing each outcome separately ignores the fact that associations might be similar across outcomes.
In this paper we propose a general approach to analyzing environmental mixtures that applies in single time point or time-varying exposure settings, which additionally borrows information across multiple outcomes to improve statistical power. In the time-varying exposure setting, we present a novel multivariate distributed lag non-linear model (MV-DLNM) and propose a strategy for borrowing strength across all exposure-outcome pairs by incorporating priors that induce clustering. In particular, we propose a co-clustering prior distribution that encourages similarity in both exposure-response curves as well as in time-lag profiles. Outside the time-varying exposure setting, the method applies immediately to other exposure structures (like exposures measured via multiple biomarkers) as well as settings with entirely unstructured exposures—for example borrowing strength across exposures in additive models or across outcomes in single index models. We then extend the proposed approach to fitting multiple index models (MIMs): with a minor modification, we show how the method can be used to fit MIMs when the true index structure—the number of indices and their composition—is unknown a priori. MIMs without known index groupings can be difficult to interpret, however, and we propose exposure importance measures that quantify the strength of each component’s contribution to the mixture.
2 Problem setup and modeling framework
For the $i^{th}$ study unit ($i=1,\dots,n$), suppose we observe a $K$-dimensional outcome vector $\mathbf{y}_{i}=(y_{i1},\dots,y_{iK})^{T}$, $P$ exposures at $L$ times: ${\mathbf{x}}_{i1},\dots,{\mathbf{x}}_{iP}$ where ${\mathbf{x}}_{ij}=({x}_{ij1},\cdots,{x}_{ijL})^{T}$, as well as a vector of other covariates $\bm{z}_{i}$. The goal is to quantify possibly non-linear associations between each outcome $y_{ik}$ and each time-varying exposure $\bm{x}_{ij}$ by fitting the following multivariate distributed lag non-linear model (MV-DLNM):
$y_{i1}$ $=f_{11}({\mathbf{x}}_{i1}^{T}{\bm{\omega}}_{11})+\cdots+f_{1P}({\mathbf{x}}_{iP}^{T}{\bm{\omega}}_{1P})+\bm{z}_{i}^{T}\bm{\beta}_{Z1}+\xi\sigma_{1}u_{i}+\epsilon_{i1}$ (1)
[…]
$y_{iK}$ $=f_{K1}({\mathbf{x}}_{i1}^{T}{\bm{\omega}}_{K1})+\cdots+f_{kj}({\mathbf{x}}_{iP}^{T}{\bm{\omega}}_{kj})+\bm{z}_{i}^{T}\bm{\beta}_{ZK}+\xi\sigma_{k}u_{i}+\epsilon_{iK}.$
Here $f_{kj}(\cdot)$ is an unknown smooth exposure response curve for outcome $k$ and exposure $j$, ${\bm{\omega}}_{kj}=({\omega}_{kj1},\dots,{\omega}_{kjL})^{T}$ is a vector of unknown time-lag weights, and $\bm{\beta}_{Zk}$ is a vector of coefficients corresponding to covariates $\bm{z}_{i}$. We let $\epsilon_{ik}$ represent independent Gaussian errors, $\epsilon_{ik}\overset{iid}{\sim}N(0,\sigma^{2})$ and allow for dependence across the same unit’s outcomes via shared random effect $u_{i}\overset{iid}{\sim}N(0,1)$, scaled by the coefficient $\xi$. To identify both $f_{kj}(\cdot)$ and $\bm{\omega}_{kj}$, constraints on the weights’ magnitude (${\bm{\omega}}_{kj}^{T}{\bm{\omega}}_{kj}=1$) and sign (e.g., ${{\omega}}_{kj1}>0$) are typically imposed.
2.1 Semiparametric modeling incorporating smoothness
The target parameters of interest are the exposure-response function $f_{kj}(\cdot)$ and the lag weights ${\bm{\omega}}_{k}=({\omega}_{k1},\dots,{\omega}_{kL})^{T}$. The exposure-response function quantifies the potentially non-linear association between the $k^{th}$ outcome and the $j^{th}$ exposure, while the lag weights quantify the relative strength of an exposure’s association at each of the $L$ time periods. Importantly, the model specification allows each of the $K\times P$ outcome-exposure pairs to have (i) different non-linear exposure-response relationships $f_{kj}(\cdot)$, and (ii) different lag profiles captured by ${\bm{\omega}}_{kj}$. To facilitate estimation and improve efficiency, we incorporate smoothness conditions on both of the target parameters separately.
Smoothness over time: Exposure measurements over time are typically highly multicollinear, making estimation of the weights $\bm{\omega}_{kj}$ challenging. To improve stability, we can encourage smoothness in the weights $\bm{\omega}_{kj}$ over time via Gaussian priors with precision matrix $\lambda^{\omega}\mathbf{D}_{r}^{T}\mathbf{D}_{r}$, where $\mathbf{D}_{r}$ is a known difference matrix of order $r$ (Gasparrini et al.,, 2017). A prior with this covariance structure effectively penalizes or shrinks differences between coefficients at similar lags, enforcing our prior belief that effects should not change drastically over small time periods. The degree of shrinkage is driven by $\lambda^{\omega}$, an unknown parameter to be estimated.
While this prior encourages smoothness over time, it leaves the magnitude of ${\bm{\omega}}_{kj}$ completely unconstrained. To address this, we propose an analogous Fisher-Bingham prior (Kent,, 1982), which corresponds to a Gaussian distribution constrained to the unit $L$-sphere, thereby enforcing the constraint ${\bm{\omega}}_{kj}^{T}{\bm{\omega}}_{kj}=1$. Incorporating the same smoothness inducing covariance structure yields the prior $\bm{\omega}_{kj}\sim FB(\mathbf{0},\lambda^{\omega}\mathbf{D}_{r}^{T}\mathbf{D}_{r})$. This addresses the magnitude constraint required for $\bm{\omega}_{kj}$. To further satisfy the sign constraint, we can restrict this prior to the half hypersphere where $\omega_{kjL}\geq 0$. The constraints on the parameter space complicate Markov chain Monte Carlo (MCMC) sampling; in Section 4 we outline two approaches to sampling these parameters.
Sampling $\bm{\omega}_{kj}$ can be computationally burdensome when $L$ is large, however. To facilitate sampling, we can incorporate a dimension reduction approach similar to the one explored for DLMs in Wilson et al., (2022). We approximate ${\bm{\omega}}_{kj}\approx\mathbf{\Psi}{\bm{\theta}}_{kj}$, where $\mathbf{\Psi}$ is a (known) matrix whose columns define an orthonormal basis expansion, and ${\bm{\theta}}_{kj}$ is an $m$-dimensional vector of unknown coefficients, where ${\bm{\theta}}_{kj}$ inherits the magnitude constraint (${\bm{\theta}}_{kj}^{T}{\bm{\theta}}_{kj}=1$) by orthonormality. Smaller values of $m$ ($m\leq L$) lead to worse approximations, but easier posterior sampling. The covariance matrix that shrinks towards smoother functions is inherited as well, yielding the proposed prior $\bm{\theta}_{kj}\sim FB(\mathbf{0},\lambda^{\omega}\mathbf{\Psi}^{T}\mathbf{D}_{r}^{T}\mathbf{D}_{r}\mathbf{\Psi})$, and we can again impose a sign constraint ${{\theta}}_{kjL}>0$, restricting to the half-sphere. In sum, we apply a known linear transformation to ${\mathbf{x}}_{ij}$ and re-parameterize ${\mathbf{x}}_{ij}^{T}{\bm{\omega}}_{kj}\approx\tilde{\bm{x}}_{ij}^{T}{\bm{\theta}}_{kj}$. This dimension reduction is not required, and we could instead perform inference directly on $\bm{\omega}_{kj}$. This approximation simply serves to improve sampling; in Section 5 we explore the degree to which it affects inference. Additionally, if ${\mathbf{x}}_{ij}$ does not represent an exposure measured over time, but instead some other structure—say, an exposure measured via several biomarkers—we simply set $\mathbf{x}_{ij}=\tilde{\mathbf{x}}_{ij}$ and ${\bm{\omega}}_{kj}={\bm{\theta}}_{kj}$, and adopt a generic ridge penalty prior on ${\bm{\omega}}_{kj}$.
Smoothness of exposure-response functions: We encourage smoothness of $f_{kj}(\cdot)$ as well, as we expect small changes in exposure levels to lead to small changes in outcomes. We approximate $f_{kj}(\cdot)$ via B-splines: $f_{kj}(\tilde{{\mathbf{x}}}_{ij}^{T}{\bm{\theta}}_{kj})\approx\mathbf{b}_{ij,\bm{\theta}_{kj}}^{T}\bm{\beta}_{kj}$ where $\mathbf{b}_{ij,\bm{\theta}_{kj}}=(b_{1}(\tilde{\mathbf{x}}_{ij}^{T}\bm{\theta}_{kj}),\dots,b_{d}(\tilde{\mathbf{x}}_{ij}^{T}\bm{\theta}_{kj}))^{T}$, $b_{1}(\cdot),\dots,b_{d}(\cdot)$ are known basis functions, and $\bm{\beta}_{kj}=(\beta_{kj1},\dots,\beta_{kjd})^{T}$ is a vector of unknown coefficients. For identifiability, these basis representations are subject to centering constraints (Wood,, 2017), and we include outcome-specific intercepts $\beta_{k0}$, $k=1,\dots,K$. To encourage smoothness, we specify Gaussian priors on $\bm{\beta}_{kj}$ with precision matrix $\lambda^{\beta}\bm{\Sigma}^{-1}_{0}$, where $\bm{\Sigma}_{0}$ is a matrix of integrated second derivatives of basis functions that penalize overly rough functions (Wood,, 2017), and $\lambda^{\beta}$ is a smoothing parameter to be estimated.
2.2 Prior distributions inducing clustering
The proposed model allows all $K\times P$ exposure-outcome pairs to have distinct time-lag profiles ($\bm{\omega}_{kj}\neq\bm{\omega}_{k^{\prime}j^{\prime}}$) and separate functional relationships ($\bm{\beta}_{kj}\neq\bm{\beta}_{k^{\prime}j^{\prime}}$). In small samples, or when the correlation among exposures across time is high, as is common in environmental epidemiology, this may be too flexible and lead to highly variable estimates. This can negatively affect power, which is already limited in the common situation where exposure effects are weak. To rein in this flexibility, we propose to encourage similarity in both $f_{kj}(\cdot)$ and in ${\bm{\omega}}_{kj}$ across different outcomes and exposures. This is a reasonable assumption in practice as many exposures have similar exposure-response curves, and the same exposure may have similar associations with different outcomes.
A natural way to borrow information would be to adopt independent DP priors. Specifically, we can set $\bm{\beta}_{kj}\sim G_{\beta}$ where $G_{\beta}\sim DP(\alpha_{\beta},G_{0,\beta})$, and $\bm{\theta}_{kj}\sim G_{\theta}$ where $G_{\theta}\sim DP(\alpha_{\theta},G_{0,\theta})$. Here $(\alpha_{\beta},\alpha_{\theta})$ are concentration parameters, dictating the degree of clustering, and $(G_{0,\beta},G_{0,\theta})$ are centering distributions, representing the means of the prior distributions. To see how this encourages clustering and borrowing of information across exposure-outcome pairs, consider the following latent variable representation (Sethuraman,, 1994):
$\bm{\beta}_{kj}$ $=\sum_{c=1}^{\infty}I(Z^{\beta}_{kj}=c)\bm{\beta}^{*}_{c}$ $\bm{\theta}_{kj}$ $=\sum_{c=1}^{\infty}I(Z^{\theta}_{kj}=c)\bm{\theta}^{*}_{c}$
$P(Z^{\beta}_{kj}=c)$ $=\pi_{c}^{\beta}$ $P(Z^{\theta}_{kj}=c)$ $=\pi_{c}^{\theta}~{}~{}c=1,2,\dots$
$\pi^{\beta}_{c}$ $=V^{\beta}_{c}\prod_{j<c}(1-V^{\beta}_{j})$ $\pi^{\theta}_{c}$ $=V^{\theta}_{c}\prod_{j<c}(1-V^{\theta}_{j})~{}~{}c=2,3,\dots,$
$V^{\beta}_{c}$ $\sim\text{Beta}(1,\alpha_{\beta})$ $V^{\theta}_{c}$ $\sim\text{Beta}(1,\alpha_{\theta}),~{}~{}c=1,2,\dots,$
$\bm{\beta}^{*}_{c}$ $\sim G_{0,\beta}$ $\bm{\theta}^{*}_{c}$ $\sim G_{0,\theta}.$
Here $(Z^{\beta}_{kj},Z^{\theta}_{kj})$ is a discrete latent variable that indicates which cluster the exposure-outcome pair belongs for the exposure-response function and lag weights, respectively. In practice we approximate the infinite mixture with a finite mixture that is capped at $C$ clusters ($V^{\beta}_{C}=V^{\theta}_{C}=1$). It is easy to show that the prior probabilities of clustering are given by $P(\bm{\beta}_{kj}=\bm{\beta}_{k^{\prime}j^{\prime}})=(\alpha_{\beta}+1)^{-1}$ and $P(\bm{\theta}_{kj}=\bm{\theta}_{k^{\prime}j^{\prime}})=(\alpha_{\theta}+1)^{-1}$. This implies that clustering will occur for both parameters, and becomes increasingly likely with smaller values of $(\alpha_{\beta},\alpha_{\theta})$.
Although this encourages clustering separately for $\bm{\beta}_{kj}$ and $\bm{\theta}_{kj}$, it does not encourage borrowing of information between the two. In practice, exposures may behave similarly both in terms of the shape of the exposure-response relationship as well as time-lag profile. For example two pollutants may both have U-shaped effects that decay linearly over the course of a week. Further, different outcomes may be susceptible to exposures at similar times and in similar manners. With that in mind we extend the DP clustering strategy above to encourage borrowing information across the exposure-response function and the time-lag domains. Consider the same latent variable representation as above. Instead of independently specifying marginal probabilities of clustering indicators $Z_{kj}^{\beta}$ and $Z_{kj}^{\theta}$, we specify a joint probability distribution that allows for dependence between the two:
$P(Z_{kj}^{\beta}=a,Z_{kj}^{\theta}=b)$ $=\frac{(1+\rho)^{I(a=b)}{\pi}^{\beta}_{a}{\pi}^{\theta}_{b}}{\sum_{a^{\prime}=1}^{C}\sum_{b^{\prime}=1}^{C}(1+\rho)^{I(a^{\prime}=b^{\prime})}{\pi}^{\beta}_{a^{\prime}}{\pi}^{\theta}_{b^{\prime}}},~{}~{}a,b\in\{1,\dots,C\}$
Here $\pi_{a}^{\beta}$ and $\pi_{b}^{\theta}$ are assigned the same stick-breaking prior distributions as in the independent DP formulation described above, and, as before, the concentration parameters $\alpha_{\beta}$ and $\alpha_{\theta}$ control the level of clustering for $\bm{\beta}_{kj}$ and $\bm{\theta}_{kj}$. However, $\rho$ now quantifies the degree of dependence in clustering between $\bm{\beta}_{kj}$ and $\bm{\theta}_{kj}$. When $\rho=0$, the model reduces to the independent DP prior case. When $\rho\to\infty$, $P(Z_{\beta,j}=a,Z_{\theta,j}=b)=1$ when $a=b$ and 0 otherwise; that is, either both $\bm{\beta}_{kj}=\bm{\beta}_{k^{\prime}j^{\prime}}$ and $\bm{\theta}_{kj}=\bm{\theta}_{k^{\prime}j^{\prime}}$, or $\bm{\beta}_{kj}\neq\bm{\beta}_{k^{\prime}j^{\prime}}$ and $\bm{\theta}_{kj}\neq\bm{\theta}_{k^{\prime}j^{\prime}}$.
We specify the centering distributions $G_{0,\beta}$ and $G_{0,\theta}$ to be the smoothness inducing priors detailed in Section 2.1. To complete the model specification, we adopt default gamma priors on ${\alpha_{\beta},\alpha_{\theta}}$, $\rho$, and smoothness parameters ${\lambda^{\beta},\lambda^{\theta}}$. For full details, see Appendix A.
3 Extension to Multiple Index Models
When exposures are not time varying, the proposed approach extends immediately to a (multivariate) single index model. Let ${\mathbf{x}}_{i}$ now represent an $L$-vector of different exposures measured for the $i^{th}$ observation. (Note: we represent the number of exposures by $L$ here rather than $P$ as in the previous section.) The multivariate single index model is:
$y_{ik}$ $=f_{k}({\mathbf{x}}_{i}^{T}{\bm{\omega}}_{k})+\bm{z}_{i}^{T}\bm{\beta}_{Zk}+\xi\sigma_{k}u_{i}+\epsilon_{ik},$
for $k=1,\dots,K$. In this model, ${\mathbf{x}}_{i}^{T}{\bm{\omega}}_{k}$ represents a multipollutant index, $f_{k}(\cdot)$ characterizes the shape of the association between the $k^{th}$ outcome and the multipollutant index, and ${{\omega}}_{kj}$ quantifies the relative strength of association between the $k^{th}$ outcome and $j^{th}$ exposure. As before, we may borrow strength across different outcomes by encouraging similarity both in ${\bm{\omega}}_{k}$ and $f_{k}(\cdot)$ via the cluster inducing priors of Section 2.2. Inference follows as above, but without a penalty to encourage smoothness across weights ${\omega}_{kj}$, $j=1,\dots,L$.
This can be overly restrictive, requiring that the association between each exposure and the $k^{th}$ outcome have the same shape, differing only in magnitude. Alternatively, multiple index models (MIMs) have been used for environmental mixtures by partitioning the exposure vector ${\mathbf{x}}_{i}$ (McGee et al.,, 2021), but this requires knowing the true index structure a priori. Instead, we propose to extend this strategy to a (multivariate) MIM of maximum order $P$:
$y_{ik}$ $=\sum_{j=1}^{P}f_{kj}({\mathbf{x}}_{i}^{T}{\bm{\omega}}_{kj})+\bm{z}_{i}^{T}\bm{\beta}_{Zk}+\xi\sigma_{k}u_{i}+\epsilon_{ik},~{}~{}k=1,\dots,K.$ (2)
This is nearly identical to model (1), but instead of summing over exposures, the summation is over over multipollutant indices that share common exposures $\mathbf{x}_{i}$. Applying the priors proposed in Section 2.2 now encourages clustering across outcomes and indices rather than exposures. This has two key consequences. First, it again improves efficiency by borrowing strength across outcomes. Second, clustering across indices facilitates collapsing to a lower order ($<P$) index model. For example if $P$=2, then when $Z^{\theta}_{k1}=Z^{\theta}_{k2}$ and $Z^{\beta}_{k1}=Z^{\beta}_{k2}$, model (2) reduces to a single index model—without assuming this a priori. This marks a major advantage over standard MIMs, as we need only specify the maximum order $P$, rather than the true index structure (i.e., the number of indices and their composition).
3.1 Exposure Importance Measures
In the time-varying exposure setting of Section 2, weights $\bm{\omega}_{kj}$ have a natural interpretation as the strength of association at different lags. In a SIM, the weights have the interpretation of being the strength of association between a particular exposure and outcome. In a MIM with unknown index structure, however, $\bm{\omega}_{kj}$ does not have a natural interpretation, and it is difficult to infer which exposures are driving the mixture association. One option is to investigate component-wise exposure response curves when holding all other exposures fixed, but this can be misleading when exposures are highly correlated. To facilitate interpretation, we propose alternative exposure importance measures that can be used for complex mixture models like the proposed MIM.
Rewriting $f^{*}_{k}(\bm{x})=\sum_{j=1}^{P}f_{kj}({\mathbf{x}}^{T}{\bm{\omega}}_{kj})$ in (2) and omitting the subscript $i$ for simplicity, we first define the overall signal-to-noise ratio for the $k^{th}$ outcome as $\text{SNR}_{k}=\frac{Var[f^{*}_{k}(\bm{X})]}{Var(\xi\sigma_{k}u+\epsilon_{k})},$ where the numerator represents the variability explained by the mixture (the signal), and the denominator represents unexplained variability (the noise). This quantifies the overall strength of the association between outcome $k$ and the exposure mixture, but does not provide information about which exposures are driving that association. By further decomposing the numerator, we construct variable importance measures for each exposure-outcome pair:
$\Psi_{kj}=1-\frac{Var_{\bm{X}_{-j}}\bigg{\{}E_{x_{j}\mid\bm{x}_{-j}}\Big{[}f^{*}_{k}(x_{j},\bm{x}_{-j})\Big{]}\bigg{\}}}{Var[f^{*}_{k}(\bm{x})]},$ (3)
where $\bm{x}_{-j}$ represents all exposures except the $j^{th}$ exposure. $\Psi_{kj}$ represents the proportion of the signal explained only by the $j^{th}$ exposure. Naturally, $\Psi_{kj}$ will be small if $x_{j}$ is not predictive of the outcome. But it will also be small if $x_{j}$ is highly collinear with $\bm{x}_{-j}$, as the association with $x_{j}$ can be easily captured by other components. In this way $\Psi_{kj}$ quantifies the importance of the $j^{th}$ exposure to the mixture effect on the $k^{th}$ outcome. In Appendix C we generalize this to arbitrary groups of exposures.
Computing these quantities is straightforward given posterior samples of $f_{k}^{*}(\bm{x})$. Marginal variances are calculated using the empirical distribution of the exposures. Estimating $\Psi_{kj}$ requires integrating over the conditional distribution of $x_{j}$ given the other exposures. If the dimension of $\bm{x}$ is small, we can estimate the conditional mean in (3) non-parametrically:
$\widehat{E}_{x_{j}\mid\bm{x}_{-j}}\{f^{*}_{k}(x_{j},\bm{x}_{-j})\}=\frac{\sum_{i=1}^{n}f^{*}_{k}(x_{ij},\bm{x}_{(-j)})K(\bm{x}_{i(-j)}-\bm{x}_{-j})}{\sum_{i=1}^{n}K(\bm{x}_{i(-j)}-\bm{x}_{-j})}$
where $K(\cdot)$ is an appropriate kernel. Alternatively, when $\bm{x}$ is high-dimensional, one could make parametric assumptions, such as assuming that $x_{j}$ given $\bm{x}_{-j}$ is normally distributed, and the conditional mean can be estimated using any regression approach.
The proposed $\Psi_{kj}$ has connections to a number of related variable importance metrics in the literature. It is closely related to those proposed in Shin et al., (2024), which instead target effect modification and identify variables that modify associations the most. It is also related to metrics used in nonparametric regression such as Shapley values (Shapley,, 1953) or leave-one-covariate-out (LOCO; Lei et al.,, 2018) that target the extent to which mean squared error is reduced when a covariate is omitted. One important distinction is that our proposed metric does not require fitting multiple models (e.g., omitting exposure $j$). Rather, it uses a single fit along with an estimate of the distribution of the exposures. While the distribution of the exposures can be misspecified, we view $\Psi_{kj}$ as a complementary piece of information to aid interpretation, even if the exposure distribution is imperfectly estimated.
4 Posterior sampling of index parameters
We sample from the posterior via MCMC. Using a standard Gibbs sampler, sampling from the full conditionals for most parameters is straightforward (see Appendix B for details). Sampling from the full conditional for the index weights $\bm{\theta}^{*}_{c}$, however, is complicated by its constrained parameter space. We present two approaches to sampling $\bm{\theta}^{*}_{c}$: the first transforms $\bm{\theta}^{*}_{c}$ to an unconstrained space; the second exploits a Taylor approximation to obtain a closed form for the full conditional distribution and ignores the sign constraint.
4.1 Polar Coordinate Reparameterization
The constraints on $\bm{\theta}^{*}_{c}$ makes choosing reasonable proposal distributions challenging in practice. One path forward is to absorb the constraints into the parameterization by transforming to polar coordinates. Following the approach of Park et al., (2005), we let
$\theta^{*}_{c1}$ $=\sin(\phi^{*}_{c1}),~{}~{}\theta^{*}_{c{\ell}}=\sin(\phi^{*}_{c{\ell}})\prod_{j=1}^{\ell-1}\cos(\phi^{*}_{cj}),~{}\ell=2,\dots,L-1,~{}~{}\text{and}~{}\theta^{*}_{cL}=\prod_{j=1}^{L-1}\cos(\phi^{*}_{cj}),$
where $\phi^{*}_{cj}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ for $j=1,\dots,{L-1}$, which satisfies both constraints by default. Whereas the entire vector $\bm{\theta}^{*}_{c}$ is inextricably linked by its magnitude constraint, the domains of $\phi^{*}_{cj}$ do not depend on one another, so we can draw each separately from its full conditional via grid sampling or Metropolis Hastings with scaled and shifted beta proposals.
4.2 Fast Approximate Sampling via Taylor Expansion
The approach of Section 4.1 can be slow to mix in high-dimensional settings. Alternatively, we propose the following approximate method, ignoring the sign constraint and using the Fisher-Bingham distribution on the full hypersphere for $G_{0,\theta}$. A first order Taylor expansion of $f_{kj}(\tilde{\mathbf{x}}_{ij}^{T}\bm{\theta}^{*}_{c})$ around the previous value $\tilde{\bm{\theta}}^{*}_{c}$ implies that the sum of squares is approximately
$\sum_{i=1}^{n}\left[y_{ik}-\beta_{0k}-\sum_{j=1}^{P}f_{kj}(\tilde{\mathbf{x}}_{ij}^{T}\bm{\theta}^{*}_{Z_{kj}^{\theta}})-\bm{z}_{i}^{T}\bm{\beta}_{Zk}-\xi\sigma_{k}u_{i}\right]^{2}$ $\approx[\hat{\mathbf{y}}_{k}-\hat{\mathbf{X}}^{T}\bm{\theta}^{*}_{c}]^{T}[\hat{\mathbf{y}}_{k}-\hat{\mathbf{X}}^{T}\bm{\theta}^{*}_{c}],$
where $\hat{\mathbf{y}}_{k}$ is a vector with $i^{th}$ element equal to
$y_{ik}-\beta_{0k}-\sum_{j:z^{\theta}_{kj}\neq c}f_{kj}(\tilde{\mathbf{x}}_{ij}^{T}\bm{\theta}^{*}_{Z^{\theta}_{kj}})-\sum_{j:z^{\theta}_{kj}=c}f_{kj}(\tilde{\mathbf{x}}_{ij}^{T}\tilde{\bm{\theta}}^{*}_{c})-\bm{z}_{i}^{T}\bm{\beta}_{Zk}-\xi\sigma_{k}u_{i}+\sum_{j:z^{\theta}_{kj}=c}f^{\prime}_{kj}(\tilde{\mathbf{x}}_{ij}^{T}\tilde{\bm{\theta}}^{*}_{c})\tilde{\mathbf{x}}_{ij}^{T}\tilde{\bm{\theta}}^{*}_{c},$
and $\hat{\mathbf{X}}_{k}$ is a new design matrix with $i^{th}$ row equal to: $\sum_{j:z^{\theta}_{kj}=c}f^{\prime}_{kj}(\tilde{\mathbf{x}}_{ij}^{T}\tilde{\bm{\theta}}^{*}_{c})\tilde{\mathbf{x}}_{ij}^{T}.$ Both quantities require derivatives of $f_{kj}(\cdot)$, but this is straightforward to compute, as derivatives of B-splines are B-splines. The full conditional for ${\bm{\theta}}^{*}_{c}$ is then approximately proportional to
$\text{I}({\bm{\theta}^{*}_{c}}^{T}\bm{\theta}^{*}_{c}=1)\exp(\bm{\eta}^{T}\bm{\theta}^{*}_{c}-\frac{1}{2}{\bm{\theta}^{*}_{c}}^{T}\bm{\Lambda}^{-1}\bm{\theta}^{*}_{c})$ (4)
$\text{where}~{}{\bm{\eta}}$ $=\tau_{\theta}\mathbf{1}+\sum_{k,j:z^{\theta}_{kj}=c}{\sigma^{-2}_{k}}\hat{\mathbf{X}}^{T}_{k}\hat{\mathbf{y}}_{k}~{}~{}~{}\text{and}~{}~{}~{}~{}\bm{\Lambda}^{-1}=\lambda^{\theta}\bm{\Sigma}_{\theta}+\sum_{k,j:z^{\theta}_{kj}=c}{\sigma^{-2}_{k}}\hat{\mathbf{X}}_{k}^{T}\hat{\mathbf{X}}_{k}.$
This corresponds to a Fisher Bingham distribution, $\bm{\theta}^{*}_{c}\sim FB({\bm{\eta}}^{T},-\frac{1}{2}\bm{\Lambda}^{-1})$, which can be conveniently sampled from via the simdd package in R (Kent et al.,, 2018). Despite its closed form, directly sampling from a Fisher-Bingham distribution can be computationally challenging in high dimensions. Absent the indicator function, the density in (4) is proportional to a multivariate normal distribution for $\bm{\theta}^{*}_{c}$. As such, in high-dimensional settings we propose to use an approximate update, where we first draw a random variable from a $N\left(\bm{\Lambda}{\bm{\eta}},~{}\bm{\Lambda}\right)$ distribution and then project this value to the unit sphere. Empirically we have found that this approximation leads to very similar samples from the posterior distribution of $\bm{\theta}^{*}_{c}$ and is significantly faster to compute.
5 Simulation studies
We present simulations demonstrating the effectiveness of the proposed methods in two different contexts. In Simulation A, we consider time-varying exposures and demonstrate the benefit of borrowing strength across exposures and outcomes in DLNMs. In Simulation B, we consider cross-sectional exposure measurements and try to capture complex, non-linear relationships between a mixture of exposures and outcomes using MIMs. In both cases, we average results across 250 simulated datasets.
5.1 Simulation A: Distributed lag models with multiple exposures
We first focus on settings where exposures are measured repeatedly over time prior to outcomes measured at a single time point. Throughout, we set the number of outcomes to be $K=4$, the number of exposures to be $P=5$, and the number of time points to be $L=52$. We varied the sample size $n=\{500,1500\}$. To induce autocorrelation within each exposure as well as dependence among different exposures, we generated the exposures from a vector autoregressive model. Specifically, we set the exposures at the first time point to be $\bm{X}_{1}=\bm{\epsilon}_{1}$, where $\bm{\epsilon}_{1}\sim\mathcal{MVN}(\bm{0},\bm{\Sigma})$, and then generated subsequent exposures according to:
$\bm{X}_{l}=\bm{M}\bm{X}_{l-1}+\bm{\epsilon}_{l}\ \text{for }l=2,\dots,L,\quad\quad\bm{\epsilon}_{l}\sim\mathcal{MVN}(\bm{0},\bm{\Sigma}).$
We set $\bm{M}=0.85\bm{I}_{p}$ to induce a high degree of correlation over time within exposures, and we set $\bm{\Sigma}_{kj}=0.6^{|j-k|}$ to induce dependence across exposures, as is typical in mixtures of environmental exposures. For a visualization of exposure trajectories, see Appendix D.
Let $\bm{x}_{j}$ represent the vector of $L$ measurements of the $j^{th}$ exposure. Responses were generated from a multivariate normal distribution with mean:
$E[\bm{Y}]$ $=\left[\begin{array}[]{l}f_{1}\left(\bm{x}_{1}^{T}\bm{\omega}_{1}\right)+f_{1}\left(\bm{x}_{2}^{T}\bm{\omega}_{2}\right)+f_{1}\left(\bm{x}_{3}^{T}\bm{\omega}_{3}\right)+f_{1}\left(\bm{x}_{4}^{T}\bm{\omega}_{1}\right)+f_{1}\left(\bm{x}_{5}^{T}\bm{\omega}_{1}\right)\\ f_{2}\left(\bm{x}_{1}^{T}\bm{\omega}_{1}\right)+f_{2}\left(\bm{x}_{2}^{T}\bm{\omega}_{2}\right)+f_{2}\left(\bm{x}_{3}^{T}\bm{\omega}_{3}\right)+f_{2}\left(\bm{x}_{4}^{T}\bm{\omega}_{1}\right)+f_{2}\left(\bm{x}_{5}^{T}\bm{\omega}_{1}\right)\\ f_{3}\left(\bm{x}_{1}^{T}\bm{\omega}_{1}\right)+f_{3}\left(\bm{x}_{2}^{T}\bm{\omega}_{2}\right)+f_{3}\left(\bm{x}_{3}^{T}\bm{\omega}_{3}\right)+f_{3}\left(\bm{x}_{4}^{T}\bm{\omega}_{1}\right)+f_{3}\left(\bm{x}_{5}^{T}\bm{\omega}_{1}\right)\\ f_{4}\left(\bm{x}_{1}^{T}\bm{\omega}_{1}\right)+f_{4}\left(\bm{x}_{2}^{T}\bm{\omega}_{2}\right)+f_{4}\left(\bm{x}_{3}^{T}\bm{\omega}_{3}\right)+f_{4}\left(\bm{x}_{4}^{T}\bm{\omega}_{1}\right)+f_{4}\left(\bm{x}_{5}^{T}\bm{\omega}_{1}\right)\end{array}\right],$
where $f_{1}(a)=0.04a$, $f_{2}(a)=0.24(0.3a)^{2}$, $f_{3}(a)=0.09a$, $f_{4}(a)=2\sin(0.2a)$; $\bm{\omega}_{1}$ is flat, $\bm{\omega}_{2}$ is increasing, and $\bm{\omega}_{3}$ is decreasing. The lag profiles were the same across outcomes, and the exposure response functions were the same across exposures. Ultimately, of the $K\times P$=20 different pairs, there were 12 distinct outcome-exposure relationships.
To each simulated dataset we then fit four variations of the proposed approach:
  1. 1.
    The MV-DLNM with dimension reduction as in Section 2.1 (DLNM dimension reduction)
  2. 2.
    The MV-DLNM with no dimension reduction (DLNM)
  3. 3.
    The MV-DLNM without cluster-inducing priors (DLNM no clustering)
  4. 4.
    The DLNM applied to each outcome separately (DLNM separate). This allows for clustering across exposures but not across outcomes.
We report mean squared error (MSE) for the posterior mean of $f_{kj}(\cdot)$ and $\bm{\omega}_{kj}$, as well as 90% credible interval coverage, averaged over all parameters and datasets. For the proposed MV-DLNM, we also report pairwise posterior clustering probabilities—$P(\bm{\beta}_{kj}=\bm{\beta}_{k^{\prime}j^{\prime}}|\bm{y})$ and $P(\bm{\omega}_{kj}=\bm{\omega}_{k^{\prime}j^{\prime}}|\bm{y})$—averaged across datasets.
Figure 1 displays average pairwise clustering probabilities. Overall the proposed framework was able to capture similarity in both $f_{kj}(\cdot)$ and $\bm{\omega}_{kj}$. There was a high degree of clustering for the $\bm{\beta}_{kj}$ parameters within outcomes, as characterized by the block diagonal structure in the top left panel, and a low degree of clustering elsewhere. One exception was that there was a moderate degree of clustering between outcomes 1 and 3 when $n=500$; the proposed clustering mechanism was capturing the fact that both had the same shape of exposure response relationship (both linear), despite being of different strengths. This improved with higher sample size, with the proposed approach erroneously clustering outcomes 1 and 3 much less frequently (bottom left panel; Figure 1). Pairwise clustering probabilities were somewhat lower for $\bm{\omega}_{kj}$ (top right panel; Figure 1), as the lag weights were only weakly identified due to high autocorrelation (and exacerbated by correlation among different exposures). Again this issue improved with higher sample size (bottom right panel; Figure 1), with pairwise similarity in lag profiles being captured much more often.
Figure 1: Posterior probabilities of pairwise clustering for both sets of model parameters, averaged across simulated datasets when $n=500$ (top) and $n=1500$ (bottom).
In Figure 2 we report MSE for $\bm{\omega}_{kj}$ when $n=500$. For brevity, we highlight results for the first two outcomes and first three exposures (with other results being similar). Clustering generally improved estimation of $\bm{\omega}_{kj}$: models with more clustering (DLNM, DLNM dimension reduction) had lower MSE than others. The benefits were largest for exposure 1 as it shared lag profiles with exposures 4 and 5, hence allowed for more borrowing of information. We also observe a tradeoff: whereas the DLNM with dimension reduction dramatically improved computation time, it yielded slightly higher MSE than without. We found similar results for estimation of $f_{kj}(\cdot)$, with some minor differences (see Appendix D). While clustering was generally beneficial for estimation of $f_{kj}(\cdot)$, clustering only across exposures (DLNM separate) outperformed full clustering for outcomes 1 and 3. As mentioned above, this is because $f_{j1}(\cdot)$ and $f_{j^{\prime}3}$ were both linear but with different slopes, and the DLNM occasionally incorrectly clustered them together—but again this improved substantially with higher sample size (Appendix D.3).
Figure 2: Mean squared error for estimating $\bm{\omega}_{kj}$ across all estimators. MSE values are scaled so that DLNM has an MSE of 1 in all scenarios.
Figure D9 in Appendix D shows that all four models yielded near nominal 90% coverage for 16 of the 20 exposure-outcome pairs, though coverage was somewhat lower for $(k,j)=$ (1,2) and (1,3) (and analogously for outcome 3). Lag weights $\bm{\omega}_{kj}$ for these pairs were commonly estimated to be flat as a result of being incorrectly clustered with other exposures for outcome 1. This shows the bias-variance trade-off inherent to clustering: MSE was generally lower for the clustering models, but this came from a significant reduction in variance and a slight increase in bias (which can affect coverage). As mentioned above, clustering improved in larger samples, and this led to better coverage. We note that this problem of under-coverage for these lag weights occurred even in the absence of clustering (albeit less so) due to the weak identifiability of parameters.
5.2 Simulation B: Single time point mixture modeling
Now we shift focus to settings where the exposures are only observed at a single time point, but index models can be used to model non-linear effects of exposures on the outcomes.
We generated datasets with $K$=4 outcomes and $P$=10 exposures and varied the sample size $n\in\{200,600,1000\}$. The exposures $\bm{x}$ were generated from a multivariate normal distribution with a moderate autoregressive correlation structure ($\rho$=0.5). Exposure-outcome associations were weak to moderate, with signal to noise ratios between 0.07 and 0.37, and we considered three scenarios that vary in their favorability to MIMs:
Scenario 1: $E[\bm{Y}]=\left[\begin{array}[]{c}0.5f_{1}(\bm{x})\\ 0.5f_{1}(\bm{x})\\ 0.5f_{2}(\bm{x})\\ 0.3f_{2}(\bm{x})\end{array}\right]$ $f_{1}(\bm{x})=\bm{x}^{T}\bm{\alpha}_{1},$
$f_{2}(\bm{x})=\bm{x}^{T}\bm{\alpha}_{1}+(\bm{x}^{T}\bm{\alpha}_{2})^{2},$
$\bm{\alpha}_{1}=(0.1,0.1,0.2,0.2,0.25,0.1,0.05,0.08,0.3,0.1)^{T}$
$\bm{\alpha}_{2}=(0.3,0.05,0.1,0.2,0.1,0.2,\text{-}0.2,\text{-}0.2,0.1,0.2)^{T}.$
Scenario 2: $E[\bm{Y}]=\left[\begin{array}[]{c}0.25f_{1}(\bm{x})\\ 0.25f_{1}(\bm{x})\\ 0.3f_{2}(\bm{x})\\ 0.3f_{2}(\bm{x})\end{array}\right]$ $f_{1}(\bm{x})=e^{0.5x_{1}}+\dfrac{1.5e^{1.2x_{2}}}{1+e^{1.2x_{2}}}-0.5x_{3}^{2},$
$f_{2}(\bm{x})=x_{9}-0.75x_{10}^{2}.$
Scenario 3: $E[\bm{Y}]=\left[\begin{array}[]{c}f_{1}(\bm{x})\\ f_{1}(\bm{x})\\ \sqrt{0.5}f_{1}(\bm{x})\\ \sqrt{0.5}f_{1}(\bm{x})\end{array}\right]$ $f_{1}(\bm{x})=(\bm{x}^{T}\bm{\alpha}_{1})\times(\bm{x}^{T}\bm{\alpha}_{2}),$
$\bm{\alpha}_{1},$ $\bm{\alpha}_{2}$ are as above.
To each dataset we fit four different mixtures approaches:
  1. 1.
    The proposed MIM applied separately to each outcome (MIM separate)
  2. 2.
    The proposed multivariate MIM, which clusters across outcomes (MIM clustering)
  3. 3.
    Bayesian kernel machine regression (BKMR, Bobb et al.,, 2015), a widely used approach in environmental epidemiology that models potentially non-linear and non-additive relationships using a Gaussian process that also performs variable selection on the exposures.
  4. 4.
    Weighted quantile sum regression (WQS, Carrico et al.,, 2015), a popular mixtures approach that fits a linear index model on exposure quantiles.
We also computed exposure importance metrics $\Psi_{kj}$ proposed in Section 3.1 for the MIM. To estimate the distribution of $x_{j}$ given $\bm{x}_{-j}$ we assumed normality and estimated the mean of $x_{j}$ given $\bm{x}_{-j}$ via random forests to allow for flexible associations between exposures. We report MSE for the posterior mean of the exposure-response surface averaged over 200 randomly chosen locations, and we also report the average exposure importance metrics.
We report results for MSE in Figure 3. In Scenario 1, the true exposure-response relationships fit into the index framework, and the MIM approaches naturally performed best. By contrast, Scenario 2 is favorable to BKMR, as the regression surfaces are sparse (70-80% of effects are null), and each non-null effect has a different functional form. As such, BKMR performed best here in small sample sizes. Surprisingly, however, the proposed MIM approaches began to outperform BKMR in larger samples, as BMKR tended to incorrectly exclude important exposures in some datasets. Scenario 3, characterized by many pairwise interactions, is not particularly favorable to either approach. Here the MIMs—despite having an additive structure—greatly outperformed the other approaches. Across all scenarios and sample sizes, MIM clustering successfully exploited shared structure across outcomes, and outperformed MIM separate (which did not borrow information across outcomes). Overall, the proposed approach fared well in a range of settings: under sparsity, it performed reasonably well, and when many exposures had small effects, it greatly outperformed competitors.
Figure 3: Mean squared error for estimating $f^{*}(\bm{x})$ at 200 randomly chosen locations for $\bm{x}$ across all estimators, sample sizes, and scenarios. MSE values are scaled so that MIM clustering has an MSE of 1 in all scenarios.
In Figure 4, we see that estimation of exposure importance metrics is difficult in small sample sizes. In particular when $n=200$ we observe substantial overestimation in the highly sparse Scenario 2, in part because when $\Psi_{kj}=0$ it can never be underestimated (as $\Psi_{kj}\geq 0$ by definition). Ultimately, as sample size increased, estimation improved significantly in all three scenarios. This shows that the proposed exposure metrics can be reasonably estimated in order to provide additional interpretability to the proposed MIM framework.
Figure 4: Exposure importance metrics for the proposed approach across all scenarios and sample sizes.
6 NMMAPS Data Analysis
The National Morbidity, Mortality and Air Pollution Study (NMMAPS) collected time series data on daily mortality, air pollution and meteorological variables in Chicago between 1987 and 2000. We observe daily cardiovascular deaths, respiratory deaths and other non-accidental deaths, as well as median PM10 and O3 measurements. Full study details can be found in Samet et al., 2000a ; Samet et al., 2000b . Mortality outcomes were log transformed and standardized. Exposures were standardized, and we excluded one two week period in 1988 with a PM10 measurement 15 standard deviations above the mean, leaving a dataset of $n=5,100$ days.
To estimate the associations between exposure and the three outcomes, we fit the proposed MV-DLNM with $K=3$ and $P=2$, with a maximum time-lag of $L=14$ days. Following Chen et al., (2019), we adjusted for day of week (categorical), seasonality (via cyclic basis splines for month, dim=4), year (via thin plate splines, dim=6), and mean daily temperature and dew point temperature both per day (via thin plate splines, dim=6) as well as averaged over the previous three days (via thin plate splines, dim=3). We allowed for different covariate associations for each outcome ($\bm{\beta}_{Z1}\neq\bm{\beta}_{Z2}\neq\bm{\beta}_{Z3}$). We fit the MV-DLNM with and without clustering to investigate the effect of borrowing strength across outcomes and exposures.
Directly comparing exposure-response curves $f_{kj}(\cdot)$ across different models can be non-intuitive, since they are functions of different estimates of $\bm{\omega}_{kj}$. Instead, similar to Chen et al., (2019), we report mean outcome differences corresponding to an increase in a pollutant from 1SD below its mean to 1SD above at each lag $l$, holding exposures at all other lags constant. We highlight results for PM10; associations for ozone are all close to null with high uncertainty (see Appendix E). Results for PM10 without clustering are reported in the left hand column of Figure 5. Associations were small across all three outcomes, because these contrasts correspond to changes in a single day exposure while holding the other 13 days’ exposures constant. Estimates varied somewhat across different outcomes. For respiratory deaths, the strongest associations occurred at lag-0, and these decayed smoothly to the null after two weeks. Associations were also strongest at lag-0 for other non-accidental deaths, although by the end of the lag window associations had the opposite sign. By contrast, the association with cardiovascular deaths was null at lag-0 but peaked at lags 7–9.
Figure 5: Mean difference in outcomes increasing PM10 from 1SD below the mean to 1SD above at each lag (holding all else constant). Three rows correspond to (log) cardiovascular, respiratory and other (non-accidental) deaths. Left column is without clustering; right column is with clustering.
Estimating $K\times P=6$ non-parametric functions $f_{kj}(\cdot)$, which are themselves functions of $6$ different unknown weight vectors $\bm{\omega}_{kj}$ of length $L=14$ makes identifying small effects challenging, so we hypothesized that the proposed clustering strategy would improve efficiency. In Figure 6 we report pairwise posterior clustering probabilities for $f_{kj}(\cdot)$ and $\bm{\omega}_{kj}$. The degree of clustering for $f_{kj}(\cdot)$ was extremely high, indicating similar associations for all outcome-exposure pairs except for the cardiovascular - O3 relationship. There was less clustering in the time-lag profiles, but results still indicate substantial similarity, especially for PM10. This is borne out by estimated associations in Figure 5 (right hand column), in which estimates are quite similar across different outcomes, with the strongest effects at lag-0 and decaying to null by the end of the lag window. In addition to improved stability, it also yielded narrower CIs, especially for respiratory deaths. Also note that the counter-intuitive negative association at time-lag of 13 in the model is removed once clustering is incorporated.
Figure 6: Pairwise posterior clustering probabilities for Chicago analysis. Left plot is for $\bm{\beta}$; right plot for $\bm{\theta}$
7 Discussion
In this paper we have proposed a general framework for modeling environmental mixtures and possibly multivariate outcomes. Although we highlighted the distributed lag framework for time-varying exposures, the proposed approach readily applies to other exposure structures. Rather than measuring $P$ pollutants at $L$ different times, for example, Bauer et al., (2020) analyzed five different biomarkers (in hair, blood, urine, nails, and saliva) of four different metals, for each of three different outcomes via separate BKMR models. Our approach can be applied to this setting by modeling each of the $P\times K=4\times 3$ outcome-exposure relationships via $f_{kj}(\bm{x}_{ij}^{T}\bm{\omega}_{kj})$, where $\bm{x}_{ij}$ is a vector of length $L=5$ representing the $p^{th}$ metal measured via each of the five different biomarkers, and $\bm{\omega}_{kj}$, representing the relative importance of each of these biomarker measurements to the overall association. Even with a mixture of univariate exposure measurements, the proposed framework extends additive models by encouraging similar exposure-response functions across different exposures and outcomes:
$y_{ik}$ $=f_{k1}({{x}}_{i1})+\cdots+f_{kj}({{x}}_{iP})+\bm{z}_{i}^{T}\bm{\beta}_{Zk}+\xi\sigma_{k}u_{i}+\epsilon_{ik},~{}~{}k=1,\dots,K.$
Here the weights are fixed ($\omega_{kj}=1$), but we can still improve performance by encouraging similarity in the functions $f_{kj}(\cdot)$. Naturally, we expect the greatest gains when $P$$\times$$K$ is large, but the proposed approach may still offer a benefit even if the outcome is univariate ($K=1$) by borrowing strength across a large number $P>1$ of exposures. Similarly, with a univariate exposure ($P=1$), the proposed approach can borrow strength across $K>1$ outcomes.
Index models offer another approach to improving power in mixtures analyses by imposing more structure than fully non-parametric approaches like BKMR. In contrast to more flexible approaches, MIMs model the exposure-response surface as a function of $M<P$ linear indices, thus reducing the dimensionality of the estimation task. When the true number of indices—or their correct compositions—is unknown, however, they can yield bias (McGee et al.,, 2021). As an extension of our development, we have shown how to fit (possibly multivariate) MIMs that are robust to misspecification of the true index structure.
Software and Data Availability
Data analyzed in Section 6 are available in the R package dlnm Gasparrini, (2011). R code for running the proposed methods are also available at github.com/glenmcgee/MVmixture.
Acknowledgements
We acknowledge the support of the Natural Sciences and Engineering Research Council of Canada (NSERC) [RGPIN-2022-03068, DGECR-2022-00433]. Research described in this article was conducted under contract to the Health Effects Institute (HEI), an organization jointly funded by the United States Environmental Protection Agency (EPA) (Assistance Award No. CR-83590201) and certain motor vehicle and engine manufacturers. The contents of this article do not necessarily reflect the views of HEI, or its sponsors, nor do they necessarily reflect the views and policies of the EPA or motor vehicle and engine manufacturers.
APPENDIX
Appendix A Model Specification
$y_{i1}$ $\approx\beta_{10}+\mathbf{b}_{i1,\bm{\theta}_{11}}^{T}\bm{\beta}_{11}+\cdots+\mathbf{b}_{iP,\bm{\theta}_{1P}}^{T}\bm{\beta}_{1P}+\bm{z}_{i}^{T}\bm{\beta}_{Z1}+\xi\sigma_{1}u_{i}+\epsilon_{i1}$
$\vdots$
$y_{iK}$ $\approx\beta_{K0}+\mathbf{b}_{i1,\bm{\theta}_{K1}}^{T}\bm{\beta}_{K1}+\cdots+\mathbf{b}_{iP,\bm{\theta}_{kj}}^{T}\bm{\beta}_{kj}+\bm{z}_{i}^{T}\bm{\beta}_{ZK}+\xi\sigma_{K}u_{i}+\epsilon_{iK}$
$\epsilon_{ik}$ $\overset{iid}{\sim}N(0,\sigma^{2})$
$u_{ik}$ $\overset{iid}{\sim}N(0,1)$
Likelihood and Parameterization
$\mathcal{L}$ $=\prod_{k=1}^{K}\prod_{i=1}^{n}f(y_{ik};\bm{\beta}_{k1},\dots,\bm{\beta}_{kj},\bm{\theta}_{k1},\dots,\bm{\theta}_{kj},u_{i})$
[…]
$\implies\mathcal{L}$ $=\prod_{k=1}^{K}\left(\prod_{a_{1}=1}^{C}\cdots\prod_{a_{p}=1}^{C}\prod_{b_{1}=1}^{C}\cdots\prod_{b_{p}=1}^{C}\left[\prod_{i=1}^{n}f(y_{ik};\bm{\beta}^{*}_{a_{1}},\dots,\bm{\beta}^{*}_{a_{p}},\bm{\theta}^{*}_{b_{1}},\dots,\bm{\theta}^{*}_{b_{p}},u_{i})\right]^{\prod_{j=1}^{p}I(Z^{\beta}_{kj}=a_{j})\prod_{j^{\prime}=1}^{p}I(Z^{\theta}_{kj^{\prime}}=b_{j})}\right)$
or $=\prod_{k=1}^{K}\left[\prod_{i=1}^{n}f(y_{ik};\bm{\beta}^{*}_{Z^{\beta}_{k1}},\dots,\bm{\beta}^{*}_{Z^{\beta}_{kj}},\bm{\theta}^{*}_{Z^{\theta}_{k1}},\dots,\bm{\theta}^{*}_{Z^{\theta}_{kj}},u_{i})\right]$
Priors
$P(Z^{\beta}_{kj}=a,Z^{\theta}_{kj}=b)$ $=\pi^{*}_{a,b}=\frac{\pi_{a,b}}{\sum_{a^{\prime}=1}^{C}\sum_{b^{\prime}=1}^{C}\pi_{a^{\prime},b^{\prime}}},~{}~{}a,b\in\{1,\dots,C\}$
$\pi_{a,b}$ $=(1+\rho)^{I(a=b)}{\pi}^{\beta}_{a\cdot}{\pi}^{\theta}_{\cdot b}$
$\pi^{\beta}_{1}$ $=V^{\beta}_{1}\qquad\qquad\qquad~{}~{}\pi^{\theta}_{1}=V^{\theta}_{1}$
$\pi^{\beta}_{c}$ $=V^{\beta}_{c}\prod_{j<c}(1-V^{\beta}_{j})\qquad\pi^{\theta}_{c}=V^{\theta}_{c}\prod_{j<c}(1-V^{\theta}_{j})~{}~{}c=2,\dots,C$
$V^{\beta}_{c}$ $\sim Beta(1,\alpha_{\beta})\qquad\qquad V^{\theta}_{c}\sim Beta(1,\alpha_{\theta}),~{}~{}c=1,\dots,C-1$
$V^{\beta}_{C}$ $=1\qquad\qquad\qquad\qquad V^{\theta}_{C}=1$
$\alpha_{\beta}$ $\sim Gamma(a_{\beta},b_{\beta})\qquad\alpha_{\theta}\sim Gamma(a_{\theta},b_{\theta})$
$\bm{\beta}^{*}_{c}$ $\sim N(\mu_{0},[\lambda^{\beta}]^{-1}\Sigma_{0})\qquad\bm{\theta}^{*}_{c}\sim FB(\tau_{\theta}\mathbf{1},~{}-\frac{1}{2}\lambda^{\theta}\Sigma_{\theta}^{-1})$
$\lambda^{\beta}$ $\sim Gamma(a_{\lambda}^{\beta},b_{\lambda}^{\beta})\qquad\lambda^{\theta}\sim Gamma(a_{\lambda}^{\theta},b_{\lambda}^{\theta})$
$\alpha_{\beta}$ $\sim Gamma(a_{\beta},b_{\beta})\qquad\alpha_{\theta}\sim Gamma(a_{\theta},b_{\theta})$
$\rho$ $\sim Gamma(a_{\rho},b_{\rho})$
$u_{i}$ $\sim N(0,1)$
$\xi$ $\sim InvGamma(a_{\xi},b_{\xi})$
$\sigma^{2}_{k}$ $\sim InvGamma(a_{\sigma},b_{\sigma})$
$P(\beta_{Zk})$ $\propto 1$
$P(\beta_{0k})$ $\propto 1$
Note we set $\tau_{\theta}=0$ if using the polar transformation method; otherwise computing the appropriate normalizing constant necessary for updating $\lambda^{\theta}$ is challenging.
Appendix B MCMC Sampler
  1. 1.
    For $k=1,\dots,K$, $j=1,\dots,p$, sample cluster IDs:
    $P(Z^{\beta}_{kj}=a|...)$ $=\frac{\pi_{a,Z^{\theta}_{kj}}\tilde{f}^{\beta}(a)}{\sum_{c=1}^{C}\pi_{c,Z^{\theta}_{kj}}\tilde{f}^{\beta}(c)},~{}~{}a\in\{1,\dots,C\},$
    where $\text{where }\tilde{f}^{\beta}(a)$ $=\left[\prod_{i=1}^{n}f(y_{ik};\bm{\beta}^{*}_{Z^{\beta}_{k1}},\dots,\bm{\beta}^{*}_{Z^{\beta}_{k(j-1)}},\bm{\beta}^{*}_{a},\bm{\beta}^{*}_{Z^{\beta}_{k(j+1)}},\dots,\bm{\beta}^{*}_{Z^{\beta}_{kj}},\bm{\theta}^{*}_{Z^{\theta}_{k1}},\dots,\bm{\theta}^{*}_{Z^{\theta}_{kj}},u_{i})\right]$
    $P(Z^{\theta}_{kj}=b|...)$ $=\frac{\pi_{Z^{\beta}_{kj},b}\tilde{f}^{\theta}(b)}{\sum_{c=1}^{C}\pi_{Z^{\beta}_{kj},b}\tilde{f}^{\theta}(c)},~{}~{}b\in\{1,\dots,C\}$
    where $\text{where }\tilde{f}^{\theta}(b)$ $=\left[\prod_{i=1}^{n}f(y_{ik};\bm{\beta}^{*}_{Z^{\beta}_{k1}},\dots,\bm{\beta}^{*}_{Z^{\beta}_{kj}},\bm{\theta}^{*}_{Z^{\theta}_{k1}},\dots,\bm{\theta}^{*}_{Z^{\theta}_{k(j-1)}},\bm{\theta}^{*}_{b},\bm{\theta}^{*}_{Z^{\theta}_{k(j+1)}},\dots,\bm{\theta}^{*}_{Z^{\theta}_{kj}},u_{i})\right]$
  2. 2.
    For $c=1,\dots,C-1$, sample weights $V^{\beta}_{c}$ and $V^{\theta}_{c}$ via MH:
    $(V^{\beta}_{c}|...)$ $\propto(1-V^{\beta}_{c})^{\alpha_{\beta}-1}\prod_{k=1}^{K}\prod_{j=1}^{p}\pi^{*}_{z^{\beta}_{kj},z^{\theta}_{kj}}$
    $(V^{\theta}_{c}|...)$ $\propto(1-V^{\theta}_{c})^{\alpha_{\theta}-1}\prod_{k=1}^{K}\prod_{j=1}^{p}\pi^{*}_{z^{\beta}_{kj},z^{\theta}_{kj}}.$
    We can either do grid sampling or MH. As proposals we could use the posterior under independent and non-informative clustering ($\rho=0$):
    $V^{\beta}_{c}|\rho=\gamma^{x}=\gamma^{y}=0...$ $\sim Beta(1+n^{\beta}_{c},\alpha_{\beta}+\sum_{c^{\prime}=c+1}^{C}n^{\beta}_{c^{\prime}})$
    $V^{\theta}_{c}|\rho=\gamma^{x}=\gamma^{y}=0...$ $\sim Beta(1+n^{\theta}_{c},\alpha_{\theta}+\sum_{c^{\prime}=c+1}^{C}n^{\theta}_{c^{\prime}})$
    where $n^{\beta}_{c}$ is the size of the $c^{th}$ cluster for $\beta$, and analogously for ${\theta}$.
  3. 3.
    For $c=1,\dots,C$, if $\sum_{kj}Z^{\beta}_{kj}=0$, then draw $\bm{\beta}^{*}_{c}$ from the prior. Otherwise sample $\bm{\beta}^{*}_{c}$ from full conditional:
    $\bm{\beta}^{*}_{c}|...$ $\sim N(\bm{\mu}^{*},\bm{\Sigma}^{*})$
    $\bm{\mu}^{*}$ $=\bm{\Sigma}^{*}\times\left(\bm{\mu}_{0}^{T}\lambda^{\beta}\bm{\Sigma}_{0}^{-1}+\sum_{k}^{K}\sigma^{-2}_{k}\left[\sum_{j:Z^{\beta}_{kj}=c}\mathbf{B}_{j,\bm{\theta}_{kj}}^{T}\right]\left[\mathbf{y}_{k}-\beta_{0k}\mathbf{1}-\xi\sigma_{k}\mathbf{u}-\sum_{j:Z^{\beta}_{kj}\neq c}\mathbf{B}_{j,\bm{\theta}_{kj}}\bm{\beta}_{kj}\right]\right)$
    $\bm{\Sigma}^{*}$ $=\left(\lambda^{\beta}\bm{\Sigma}_{0}^{-1}+\sum_{k}^{K}\sigma^{-2}_{k}\left[\sum_{j:Z^{\beta}_{kj}=c}\mathbf{B}_{j,\bm{\theta}_{kj}}^{T}\right]\left[\sum_{j:Z^{\beta}_{kj}=c}\mathbf{B}_{j,\bm{\theta}_{kj}}\right]\right)^{-1}$
  4. 4.
    For $c=1,\dots,C$ if $\sum_{kj}Z^{\theta}_{kj}=0$, then draw $\bm{\theta}^{*}_{c}$ from the prior. Otherwise sample $\bm{\theta}^{*}_{c}$ via one of the two following methods:
    1. (a)
      MH on polar coordinates:
      $\theta^{*}_{c1}$ $=\sin(\phi^{*}_{c1})$
      $\theta^{*}_{c2}$ $=\sin(\phi^{*}_{c2})\cos(\phi^{*}_{c1})$
      $\vdots$
      $\theta^{*}_{c{L-1}}$ $=\sin(\phi^{*}_{c{L-1}})\prod_{j=1}^{L-2}\cos(\phi^{*}_{cj})$
      $\theta^{*}_{cL}$ $=\prod_{j=1}^{L-1}\cos(\phi^{*}_{cj})$
      where $\phi^{*}_{cj}\in[-\frac{\pi}{2},\frac{\pi}{2}]$ $j=1,\dots,{L-1}$.
      $(\phi^{*}_{cj}|...)$ $\propto\exp\left(\tau_{\theta}\mathbf{1}^{T}\bm{\theta}^{*}_{c}-\frac{1}{2}\lambda^{\theta}{\bm{\theta}_{c}^{*}}^{T}\Sigma_{\theta}^{-1}\bm{\theta}_{c}^{*}\right)$
      ×$\qquad\qquad\times\prod_{k=1}^{K}\left[\prod_{i=1}^{n}f(y_{ik};\bm{\beta}^{*}_{Z^{\beta}_{k1}},\dots,\bm{\beta}^{*}_{Z^{\beta}_{kj}},\bm{\theta}^{*}_{Z^{\theta}_{k1}},\dots,\bm{\theta}^{*}_{Z^{\theta}_{kj}},u_{i})\right]^{I(\sum_{j}z^{\theta}_{kj}>0)}$
      ×$\qquad\qquad\times|\cos(\phi^{*}_{cj})^{L-j}|$
      (where the last component comes from change of variable). We use scaled/shifted Beta proposals: $Beta(a_{\phi},b_{\phi})$ with the mode set to be the previous value: i.e. $b_{\omega}=((1-\left[\frac{\phi^{*}_{cj}+\frac{\pi}{2}}{\pi}\right])a_{\phi}+2\left[\frac{\phi^{*}_{cj}+\frac{\pi}{2}}{\pi}\right]-1)/\left[\frac{\phi^{*}_{cj}+\frac{\pi}{2}}{\pi}\right]$.
    2. (b)
      Taylor Approximation + Fisher Bingham
      Approximating $f_{kj}(\tilde{\mathbf{x}}_{ij}^{T}\bm{\theta}^{*}_{c})$ around the current value $\tilde{\bm{\theta}}^{*}_{c}$ yields:
      $f_{kj}(\tilde{\mathbf{x}}_{ij}^{T}\bm{\theta}^{*}_{c})$ $\approx f_{kj}(\tilde{\mathbf{x}}_{ij}^{T}\tilde{\bm{\theta}}^{*}_{c})+f^{\prime}_{kj}(\tilde{\mathbf{x}}_{ij}^{T}\tilde{\bm{\theta}}^{*}_{c})\tilde{\mathbf{x}}_{ij}^{T}(\bm{\theta}^{*}_{c}-\tilde{\bm{\theta}}^{*}_{c})\implies$
      $\sum_{i=1}^{n}\left[\left(y_{ik}-\beta_{0k}-\sum_{j:z^{\theta}_{kj}\neq c}f_{kj}(\tilde{\mathbf{x}}_{ij}^{T}\bm{\theta}_{kj})-\xi\sigma_{k}u_{i}\right)-\sum_{j:z^{\theta}_{kj}=c}f_{kj}(\tilde{\mathbf{x}}_{ij}^{T}\bm{\theta}^{*}_{c})\right]^{2}$ $\approx[\hat{\mathbf{y}}_{k}-\hat{\mathbf{X}}^{T}\bm{\theta}^{*}_{c}]^{T}[\hat{\mathbf{y}}_{k}-\hat{\mathbf{X}}^{T}\bm{\theta}^{*}_{c}]$
      where $\hat{\mathbf{y}}_{k}$ is a vector with $i^{th}$ element
      $\left(y_{ik}-\beta_{0k}-\sum_{j:z^{\theta}_{kj}\neq c}f_{kj}(\tilde{\mathbf{x}}_{ij}^{T}\bm{\theta}_{kj})-\xi\sigma_{k}u_{i}\right)-\sum_{j:z^{\theta}_{kj}=c}f_{kj}(\tilde{\mathbf{x}}_{ij}^{T}\tilde{\bm{\theta}}_{kj})+\sum_{j:z^{\theta}_{kj}=c}f^{\prime}_{kj}(\tilde{\mathbf{x}}_{ij}^{T}\tilde{\bm{\theta}}_{kj})\tilde{\mathbf{x}}_{ij}^{T}\tilde{\bm{\theta}}_{kj}$
      $\hat{\mathbf{X}}_{k}$ is a new design matrix with $i^{th}$ row equal to:
      $\sum_{j:z^{\theta}_{kj}=c}f^{\prime}_{kj}(\tilde{\mathbf{x}}_{ij}^{T}\tilde{\bm{\theta}}_{kj})\tilde{\mathbf{x}}_{ij}^{T}$
      Then we can approximate the full conditional by
      $(\bm{\theta}^{*}_{c}|...)$ $\propto I(||\bm{\theta}^{*}_{c}||=1)\exp\left(\left[\tau_{\theta}\mathbf{1}+\sum_{k,j:z^{\theta}_{kj}=c}\sigma^{-2}_{k}\hat{\mathbf{X}}^{T}_{k}\hat{\mathbf{y}}_{k}\right]^{T}\bm{\theta}^{*}_{c}\right.$
      $\qquad\qquad\qquad\qquad\qquad-\frac{1}{2}{\bm{\theta}^{*}_{c}}^{T}\left[\sum_{k,j:z^{\theta}_{kj}=c}\sigma^{-2}_{k}\hat{\mathbf{X}}_{k}^{T}\hat{\mathbf{X}}_{k}\right]\bm{\theta}^{*}_{c}-\frac{1}{2}\lambda^{\theta}{\bm{\theta}_{c}^{*}}^{T}\Sigma_{\theta}^{-1}\bm{\theta}_{c}^{*}\Biggr{)}$
      $\bm{\theta}^{*}_{c}|...$ $\sim FB\left(\bar{\bm{\mu}}=\left[\tau_{\theta}\mathbf{1}+\sum_{k,j:z^{\theta}_{kj}=c}\sigma^{-2}_{k}\hat{\mathbf{X}}^{T}_{k}\hat{\mathbf{y}}_{k}\right]^{T},~{}~{}\bar{\mathbf{A}}=-\frac{1}{2}\left[\sum_{k,j:z^{\theta}_{kj}=c}\sigma^{-2}_{k}\hat{\mathbf{X}}_{k}^{T}\hat{\mathbf{X}}_{k}\right]-\frac{1}{2}\lambda^{\theta}\Sigma_{\theta}^{-1}\right)$
      In practice sampling from this distribution can be challenging, so we approximate it by sampling from $N\left(\hat{\bm{\mu}},~{}~{}\hat{\bm{\Sigma}}\right)$ and then standardizing appropriately:
      $\hat{\bm{\mu}}$ $=\left[\left[\sum_{k,j:z^{\theta}_{kj}=c}\sigma^{-2}_{k}\hat{\mathbf{X}}_{k}^{T}\hat{\mathbf{X}}_{k}\right]+\lambda^{\theta}\Sigma_{\theta}^{-1}\right]^{-1}\left[\tau_{\theta}\mathbf{1}+\sum_{k,j:z^{\theta}_{kj}=c}\sigma^{-2}_{k}\hat{\mathbf{X}}^{T}_{k}\hat{\mathbf{y}}_{k}\right]^{T}$
      $\hat{\bm{\Sigma}}$ $=\left[\left[\sum_{k,j:z^{\theta}_{kj}=c}\sigma^{-2}_{k}\hat{\mathbf{X}}_{k}^{T}\hat{\mathbf{X}}_{k}\right]+\lambda^{\theta}\Sigma_{\theta}^{-1}\right]^{-1}$
  5. 5.
    Sample $\alpha_{\beta}$ and $\alpha_{\theta}$ from full conditionals:
    $\alpha_{\beta}|...$ $\sim Gamma\left(a_{\beta}+C-1,b_{\beta}-\sum_{c=1}^{C-1}\log(1-V^{\beta}_{c})\right)$
    $\alpha_{\theta}|...$ $\sim Gamma\left(a_{\theta}+C-1,b_{\theta}-\sum_{c=1}^{C-1}\log(1-V^{\theta}_{c})\right)$
  6. 6.
    Sample $\log(\rho)$ via random walk metropolis:
    $(\log(\rho)|...)$ $\propto Gamma(a_{\rho},b_{\rho})\times\prod_{k=1}^{K}\prod_{j=1}^{p}\pi^{*}_{z^{\beta}_{kj},z^{\theta}_{kj}}\times\rho$
    (where the last component comes from the change of variable).
  7. 7.
    Sample penalty term $\lambda^{\beta}$ from full conditional
    $\lambda^{\beta}|\dots$ $\sim Gamma\left(a_{\lambda}^{\beta}+\frac{Cd}{2},~{}~{}~{}b_{\lambda}^{\beta}+\frac{1}{2}\sum_{c=1}^{C}{\bm{\beta}_{c}^{*}}^{T}\bm{\Sigma}_{0}^{-1}{\bm{\beta}_{c}^{*}}\right)$
    where $d$ is the dimension of $\bm{\beta}^{*}_{c}$.
  8. 8.
    Sample penalty term $\log\lambda^{\theta}$ via random walk metropolis (if using DLM smoothness penalty):
    $(\log(\lambda^{\theta})|\dots)$ $\propto Gamma(a_{\lambda}^{\theta},b_{\lambda}^{\theta})\times[C_{0}(\tau_{\theta}\mathbf{1},-\frac{1}{2}\lambda^{\theta}\Sigma_{\theta}^{-1})]^{-C}\exp\left(-\frac{1}{2}\lambda^{\theta}\sum_{c=1}^{C}{\bm{\theta}_{c}^{*}}^{T}\Sigma_{\theta}^{-1}\bm{\theta}_{c}^{*}\right)\times\lambda^{\theta}$
    where $C_{0}(\cdot)$ is the normalizing constant for the FB distribution (which can be obtained via fb.saddle() in the Directional package).
  9. 9.
    For $i=1,\dots,n$ sample random effects $u_{i}$ from full conditional:
    $u_{i}|...$ $\sim N\left(\left[1+\sum_{k=1}^{K}\frac{\xi^{2}\sigma_{k}^{2}}{\sigma^{2}_{k}}\right]^{-1}\sum_{k=1}^{K}\frac{\xi\sigma_{k}}{\sigma^{2}_{k}}\left[y_{ik}-\beta_{0k}-\sum_{j=1}^{p}\mathbf{b}_{i,\bm{\theta}_{kj}}^{T}\bm{\beta}_{kj}\right],~{}\left[1+\sum_{k=1}^{K}\frac{\xi^{2}\sigma_{k}^{2}}{\sigma^{2}_{k}}\right]^{-1}\right)$
  10. 10.
    For $k=1,\dots,K$, sample random effects coefficient/covariance component $\xi$ via random walk MH on the log scale:
    $(\text{log}(\xi)|...)$ $\propto InvGamma(a_{\xi},b_{\xi})\times\exp\left(-\frac{1}{2}\sum_{k=1}^{K}\frac{1}{\sigma^{2}_{k}}\sum_{i=1}^{n}\left[y_{ik}-\beta_{0k}-\sum_{j=1}^{p}\mathbf{b}_{i,\bm{\theta}_{kj}}^{T}\bm{\beta}_{kj}-\xi\sigma_{k}u_{i}\right]^{2}\right)\times\xi$
    where the last component comes from the change of variable.
  11. 11.
    For $k=1,\dots,K$, sample error variances $\sigma^{2}_{k}$ via random walk MH on the log scale
    $(\text{log}(\sigma^{2}_{k})|...)$ $\propto InvGamma\left(a_{\sigma},b_{\sigma}\right)\times(\sigma^{2}_{k})^{-\frac{n}{2}}\exp\left(-\frac{1}{2\sigma^{2}_{k}}\sum_{i=1}^{n}\left[y_{ik}-\beta_{0k}-\sum_{j=1}^{p}\mathbf{b}_{i,\bm{\theta}_{kj}}^{T}\bm{\beta}_{kj}-\xi\sigma_{k}u_{i}\right]^{2}\right)\times\sigma^{2}_{k}$
    where the last component comes from the change of variable.
  12. 12.
    For $k=1,\dots,K$, sample intercept $\beta_{0k}$ from full conditional:
    $\beta_{0k}|...$ $\sim N\left(\frac{1}{n}\sum_{i=1}^{n}\left[y_{ik}-\sum_{j=1}^{p}\mathbf{b}_{i,\bm{\theta}_{kj}}^{T}\bm{\beta}_{kj}-\xi\sigma_{k}u_{i}\right]^{2},~{}~{}\sigma^{2}_{k}/n\right)$
  13. 13.
    For $k=1,\dots,K$ sample confounder coefficients $\bm{\beta}_{Zk}$ from full conditional:
    $\bm{\beta}_{Zk}|\cdots\sim N\left(\left[\mathbf{Z}^{T}\mathbf{Z}\right]^{-1}\mathbf{Z}^{T}\left[\textbf{y}_{\cdot k}-\beta_{0k}-\sum_{j=1}^{p}\mathbf{b}_{i,\bm{\theta}_{kj}}^{T}\bm{\beta}_{kj}-\xi\sigma_{k}\mathbf{u},\right],~{}~{}\sigma_{k}^{2}\left[\mathbf{Z}^{T},\mathbf{Z}\right]^{-1}\right)$
Appendix C Variable Importance Metrics for Arbitrary Groups of Exposures
When interest lies not in individual exposures but in an important group of exposures $\bm{x}^{\dagger}$—for example a class of highly correlated exposures—our proposed importance measure $\Psi_{kj}$ generalizes to any partition $\bm{x}^{T}=[{\bm{x}^{\dagger}}^{T},{\bm{x}^{o}}^{T}]$:
$\Psi_{k}(\bm{x}^{\dagger})=1-\frac{Var_{\bm{x}^{o}}\bigg{\{}E_{\bm{x}^{\dagger}\mid\bm{x}^{o}}\Big{[}f^{*}_{k}(\bm{x}^{\dagger},\bm{x}^{o})\Big{]}\bigg{\}}}{Var[f^{*}_{k}(\bm{x})]}.$
This now represents the importance of the group of exposures found in $\bm{x}^{\dagger}$. Computation follows similarly to the univariate case.
Appendix D Additional simulation results
D.1 Illustration of Autocorrelated Exposures for Simulation A
Here we provide additional information on the data generating process for the simulations involving distributed lag exposures in Section 5.1. Figure D7 shows three realizations (three observations) of the exposure vectors measured across time. We see that there is a moderate degree of temporal dependence due to the autoregressive structure with a large correlation coefficient. Additionally, the five exposures tend to increase (decrease) together over time due to the dependence induced by the vector autoregressive process generating the data, which imposes correlation across exposures at each point in time.
Figure D7: Exposure trajectories over time for three randomly chosen observations from one of the simulated data sets in Section 5.1.
D.2 Additional results
Here we provide additional results for the simulation studies found in Section 5.1. Specifically, we now provide 90% credible interval coverage for the $\bm{\omega}_{kj}$ parameters (Figure D9), estimates of the individual $\bm{\omega}_{kj}$ trajectories across data sets and their corresponding means (Figure D10), and mean squared error of estimation of $f_{kj}(\cdot)$ (Figure D8). Overall, the results echo those seen in Section 5.1, but provide some additional insight into clustering, and when it is beneficial. We see that the interval coverages are generally at the desired 90% level, however, for outcome 1 under both exposure 2 and 3, we see some coverage below the nominal level. We see that this issue is reduced when we do not perform clustering as the DLNM clustering estimator has credible interval closer to the nominal level. Figure D10 provides intuition for this, as we can see that in many data sets, exposures 2 and 3 are placed into the wrong cluster for $\bm{\omega}_{kj}$. This generally highlights the bias-variance trade-off that comes from our clustering algorithm. Clustering generally improves overall MSE of these parameters, which are difficult to estimate in small sample sizes, but the shrinkage across exposures and outcomes induces some bias that can lead to reduced credible interval coverages. Lastly, when estimating $f_{kj}(\cdot)$, we see similar results as for $\bm{\omega}_{kj}$, though there are certain instances where it is better to not perform clustering across outcomes, as seen in the top panel of Figure D8. This issue, along with the reduced credible intervals for certain combinations of exposures and outcomes, disappears for larger sample sizes, which we show in the following section.
Figure D8: Mean squared error for estimating $f_{kj}(\cdot)$ across all estimators. MSE is scaled so that the DLNM estimator has an MSE of 1 in all scenarios.
Figure D9: 90% credible interval coverage for estimating $\bm{\omega}_{kj}$ across all estimators.
Figure D10: Estimates of $\bm{\omega}_{kj}$ for the DLNM estimator. Individual grey lines correspond to individual data sets, while the red line represents the truth, and the black line represents the mean across data sets.
D.3 Larger sample sizes
Here we present results from the same data generating process as the distributed lag simulation study in Section 5.1, however, we increase the sample size to $n=1500$. The results can be found across Figures D11 - D15. The results are largely similar to what was seen for smaller sample sizes, with a few expected, but notable, changes. Overall, the results point to the same finding that our proposed DLNM estimator generally outperforms estimators that do not perform clustering, and our approach without dimension reduction, while computationally slower, outperforms the one with dimension reduction as expected. One key difference in these results is that the clustering for the $\bm{\omega}_{kj}$ terms is significantly improved relative to the smaller sample size setting. This can be seen in Figure D12 where there is much clearer clustering structure, or in Figure D14 where there are fewer data sets for which the estimated distributed lag curves are placed in the wrong cluster. This leads to improved credible interval coverages, as seen in Figure D13. We now see that it is also preferable (in terms of MSE) to use our DLNM model with clustering across both exposures and outcomes for estimating $f_{kj}(\cdot)$, as seen in Figure D15. As a general rule, the $\bm{\omega}_{kj}$ parameters are only weakly identified, and are very difficult to estimate from small sample sizes, because of the high degree of correlation in the exposures across time. These issues improve as the sample size increases, which leads to the improved clustering performance, and the improved results when performing inference on the $\bm{\omega}_{kj}$ parameters.
Figure D11: Mean squared error for estimating $\bm{\omega}_{kj}$ across all estimators when the sample size is $n=1500$. MSE is scaled so that the DLNM estimator has an MSE of 1 in all scenarios.
Figure D12: Posterior probabilities of pairwise clustering for both sets of model parameters, averaged across simulated datasets when the sample size is $n=1500$.
Figure D13: 90% credible interval coverage for estimating $\bm{\omega}_{kj}$ across all estimators when the sample size is $n=1500$.
Figure D14: Estimates of $\bm{\omega}_{kj}$ for the DLNM estimator when the sample size is $n=1500$. Individual grey lines correspond to individual data sets, while the red line represents the truth, and the black line represents the mean across data sets.
Figure D15: Mean squared error for estimating $f_{kj}(\cdot)$ across all estimators when the sample size is $n=1500$. MSE is scaled so that the DLNM estimator has an MSE of 1 in all scenarios.
Appendix E Additional NMMAPS Analysis Results
In Figure E16 we report mean outcome differences corresponding to an increase in ozone from 1SD below its mean to 1SD above at each lag $l$, holding exposures at all other lags constant. Results without clustering are reported on the left hand column; results with clustering are reported on the right. Overall, evidence of associations with ozone was low. Both approaches yielded constant null effects of ozone on cardiovascular deaths. Associations with respiratory deaths were strongest at lag 0 across both approaches; without clustering there was a slight increase in associations at lag 14, whereas with clustering lag effects decreased monotonically with lag. In both cases estimated associations were dwarfed by uncertainty. For other (non-accidental) deaths, the non-clustering approach found small protective effects of ozone at lag 0, although there was high uncertainty. The clustering approach reduced this coutnerintuitive result, however; associations were estimated to be constant and null across all lags.
Figure E16: Mean difference in outcomes increasing O3 from 1SD below the mean to 1SD above at each lag (holding all else constant). Three rows correspond to (log) cardiovascular, respiratory and other (non-accidental) deaths. Left column is with clustering; right column without clustering.
Estimated lag weight functions for PM10 and O3 are reported in Figures E17 E18, respectively. For identifiability, weights at lag 0 were constrained to be positive. (Note that the apparent ‘reflection’ in the top row of E17 is simply an artifact of the identifiability constraint, because lag-0 weights were estimated to be close to 0.) Overall results were similar to those above. For the non-clustered approach, the PM10-cardiovascular association exhibited U-shaped lag profiles, with strongest weights at days 7–9. The O3-cardiovascular weights were effectively constant across lags, and other lags were roughly linear. Clustering appeared to have the largest effect on the PM10-cardiovascular association, with lag profiles being a combination of U-shaped functions and linear functions. The O3-cardiovascular weights remained largely constant across lags. Other exposure-outcome pairs again had linear lag profiles, although weights at lag 13 were closer to null than in the non-clustering approach.
Figure E17: Lagged weight functions for PM10 analysis, constrained to have positive weight at lag 0. Three rows correspond to (log) cardiovascular, respiratory and other (non-accidental) deaths. Left column is without clustering; right column is with clustering.
Figure E18: Lagged weight functions for O3 analysis, constrained to have positive weight at lag 0. Three rows correspond to (log) cardiovascular, respiratory and other (non-accidental) deaths. Left column is with clustering; right column without clustering.